# Characterization of Efficient Influence Function for Off-Policy Evaluation Under Optimal Policies

### v1

@@ Abstract
Off-policy evaluation (OPE) provides a powerful framework for estimating the value of a counterfactual policy using observational data, without the need for additional experimentation. Despite recent progress in robust and efficient OPE across various settings, rigorous efficiency analysis of OPE under an estimated optimal policy remains limited. In this paper, we establish a concise characterization of the efficient influence function for the value function under optimal policy within canonical Markov decision process models. Specifically, we provide the sufficient conditions for the existence of the efficient influence function and characterize its expression. We also give the conditions under which the EIF does not exist.
@@ 1 Introduction
In this paper, we provide a simple result for the characterization of the efficient influence function (EIF) for the value function under optimal policy within canonical Markov decision process models. Specifically, we establish sufficient conditions for the existence of the EIF and derive its explicit expression. We also identify and prove conditions under which the EIF does not exist, and demonstrate that various existing methods, including the value functions in Fitted Q Iteration (FQI, for example Munos & Szepesvári, 2008; Shi et al., 2024a) and Fitted Actor-Critic (FAC, see Xu et al., 2021; Kumar et al., 2023), do achieve semiparametric efficiency without parameterization or fixing the policies.
@@ 2.2 Two Estimators under Fixed Policies
under any policy $\pi$ where $f(s_{0})$ is the density of the initial state distribution. Usually, (LABEL:Q_to_eta_eq) ensures it is sufficient to learn the cumulative rewards for any given policy of interest $\pi$, and we also define a weighted average of the $Q$-function as $V(s;\pi)=\int Q(a,s;\pi)\pi(a\mid s)\,\mathrm{d}a$.
@@ 2.3 Characterization and Issues
which implies $\pi^{*}(A\mid S):=\arg\min_{\pi\in\mathcal{P}}\Delta(\pi;A,S)=\pi^{*}(P)(A,S)$. Define the new functional
@@
From Chapter 3.3 of Tsiatis (2006), we know that $\widehat{\eta}^{*}$ from (5) is an RAL estimator with the IF $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$ and thus achieves the semiparametric efficiency. Indeed, we will show that various well-known estimators in Section 4 all satisfy this display, thus are semiparametrically efficient.
@@ 3 Existence of Efficient Influence Function and Its Expression
  • Actions: the action space $\mathcal{A}$ is finite and there exist positive constants $\underline{c}_{\pi}$ and $\overline{c}_{\pi}$ such that $\underline{c}_{\pi}\leq\inf_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}$;
@@
All conditions in Assumption 3 is commonly-used in various MDP literature including OPE as well as Q-learning (for example, see the review papers Levine et al., 2020; Uehara et al., 2022; Shi, 2025), except the assumption that $0<\underline{c}_{\pi}\leq\inf_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}<\infty$. However, this assumption is also quite mild: as one can see that if $\mathcal{S}$ is finite, then this assumption will be automatically satisfied, and such assumption ($\mathcal{S}$ is finite) is also quite commonly-used in OPE literature (see Xu et al., 2021; Shi et al., 2024b; Bian et al., 2024, for instance).
@@ 4 Examples
We present several commonly-used existing methods that achieve the semiparametric efficiency without parameterization or fixing the policies.

### v2

@@ Abstract
Off-policy evaluation (OPE) provides a powerful framework for estimating the value of a counterfactual policy using observational data, without the need for additional experimentation. Despite recent progress in robust and efficient OPE across various settings, rigorous efficiency analysis of OPE under an estimated optimal policy remains limited. In this paper, we establish a concise characterization of the efficient influence function (EIF) for the value function under optimal policy within canonical Markov decision process models. Specifically, we provide the sufficient conditions for the existence of the EIF and characterize its expression. We also give the conditions under which the EIF does not exist.
@@ 1 Introduction
In this paper, we provide a simple result for the characterization of the efficient influence function (EIF) for the value function under optimal policy within canonical Markov decision process models. Specifically, we establish sufficient conditions for the existence of the EIF and derive its explicit expression under the condition that the optimal policy is unique and deterministic. We also identify and prove conditions under which the EIF does not exist. Additionally, we demonstrate that various existing methods, including the value functions in Fitted Q Iteration (FQI, for example Munos & Szepesvári, 2008; Shi et al., 2024a) and Fitted Actor-Critic (FAC, see Xu et al., 2021; Kumar et al., 2023), do achieve semiparametric efficiency without parameterization or fixing the policies in cases where the optimal policy is unique and deterministic.
@@ 2.2 Two Estimators under Fixed Policies
under any policy $\pi$ where $f(s_{0})$ is the density of the initial state distribution. Usually, the above display ensures it is sufficient to learn the cumulative rewards for any given policy of interest $\pi$, and we also define a weighted average of the $Q$-function as $V(s;\pi)=\int Q(a,s;\pi)\pi(a\mid s)\,\mathrm{d}a$.
@@ 2.3 Characterization and Issues
which also implies $\pi^{*}(P)(A,S)=\arg\min_{\pi\in\mathcal{P}}\Delta(P)(\pi;A,S)$. Define the new functional
@@
From Chapter 3.3 of Tsiatis (2006), we know that $\widehat{\eta}^{*}$ from (5) is an RAL estimator with the IF $S^{\text{eff, nonpar}}\{\Psi^{*}(P)\}$ and thus achieves the semiparametric efficiency.
@@ 3 Existence of Efficient Influence Function and Its Expression
  • Actions: the action space $\mathcal{A}$ is finite;
  • Policies: There exist positive constants $\underline{c}_{\pi}$ and $\overline{c}_{\pi}$ such that $\underline{c}_{\pi}\leq\inf_{\pi\in\mathcal{P}}\operatorname{essinf}_{(a,s)\in\mathcal{A}\times\mathcal{S}}\pi(a\mid s)\leq\sup_{\pi\in\mathcal{P}}\|\pi\|_{\infty}\leq\overline{c}_{\pi}$;
@@
All conditions in Assumption 3 is commonly-used in various MDP literature including OPE as well as $Q$-learning (for example, see the review papers Levine et al., 2020; Uehara et al., 2022; Shi et al., 2022; Shi, 2025), except the assumptions for the policies. However, this assumption is also quite mild: as one can see that if $\mathcal{S}$ is finite, then it will be automatically satisfied, and such assumption ($\mathcal{S}$ is finite) is also quite commonly-used in OPE literature (see Xu et al., 2021; Shi et al., 2024b; Bian et al., 2024, for instance).
@@ 4 Examples
We present several commonly-used existing methods that achieve semiparametric efficiency without parameterization or fixing the policies in cases where the optimal policy is unique and deterministic.